OpenStudy (anonymous):

Can some one help with this Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n ) = f(1) + f(2) + f(n - 1), for n > 2

OpenStudy (anonymous):

@ganeshie8 can u please help mee

ganeshie8 (ganeshie8):

recursive sequence !

ganeshie8 (ganeshie8):

that means you cannot find f(5) without finding the f(4) first

ganeshie8 (ganeshie8):

and you cannot find f(4) without finding the f(3) first and so on

OpenStudy (anonymous):

so what do i do

ganeshie8 (ganeshie8):

so lets find f(3) first : f(n ) = f(1) + f(2) + f(n - 1) plugin n = 3 : f(3) = f(1) + f(2) + f(3-1) = f(1) + f(2) + f(2) = ?

OpenStudy (anonymous):

umm f(5) idk im not good at this

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

you're given the values of f(1) and f(2) already, just plugin above ^^

OpenStudy (anonymous):

huh

ganeshie8 (ganeshie8):

don't huh meh

ganeshie8 (ganeshie8):

let me know exactly what u don't understand so that it helps me explain better >.<

OpenStudy (anonymous):

im sorry i just dont understand where all these # are comming from